### v1

# On Fairness of Task Arithmetic: The Role of Task Vectors

Abstract
Model editing techniques, particularly task arithmetic using task vectors, have shown promise in efficiently modifying pre-trained models through arithmetic operations like task addition and negation. Despite computational advantages, these methods may inadvertently affect model fairness, creating risks in sensitive applications like hate speech detection. However, the fairness implications of task arithmetic remain largely unexplored, presenting a critical gap in the existing literature. We systematically examine how manipulating task vectors affects fairness metrics, including Demographic Parity and Equalized Odds. To rigorously assess these effects, we benchmark task arithmetic against full fine-tuning, a costly but widely used baseline, and Low-Rank Adaptation (LoRA), a prevalent parameter-efficient fine-tuning method. Additionally, we explore merging task vectors from models fine-tuned on demographic subgroups vulnerable to hate speech, investigating whether fairness outcomes can be controlled by adjusting task vector coefficients, potentially enabling tailored model behavior. Our results offer novel insights into the fairness implications of model editing and establish a foundation for fairness-aware and responsible model editing practices.
On Fairness of Task Arithmetic: The Role of Task Vectors
Hiroki Naganumathanks: Equal contributions. thanks: Correspondence: naganuma.hiroki@mila.quebec Université de Montréal, Mila                        Kotaro Yoshida11footnotemark: 1 Institute of Science Tokyo                        Laura Gomezjurado Gonzalez11footnotemark: 1 Stanford University
Takafumi Horie Kyoto University                        Yuji Naraki Independent Researcher                        Ryotaro Shimizu University of California San Diego, ZOZO Research
1 Introduction
As large language models (LLMs) see broader application, efficient techniques for adapting them to specific tasks become increasingly crucial. While some models have been distilled (Sanh et al., 2019; Jiao et al., 2020; Turc et al., 2020) or are relatively small (Abdin et al., 2024), task-specific fine-tuning often demands substantial computational resources, prompting the development of parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Hu et al., 2022; Ben Zaken et al., 2022; Dettmers et al., 2023).
One notable example is Low-Rank Adaptation (LoRA) (Hu et al., 2022), which updates a compact set of parameters while leaving most of the original weights untouched, thus reducing training costs. Despite the popularity of PEFT methods, they do not resolve every challenge: in high-stakes tasks with imbalanced data, LoRA and similar approaches can preserve or even amplify biases, raising concerns about fairness (Ding et al., 2024b; Sap et al., 2019).
An alternative strategy that has recently drawn attention is model editing with task vectors (Ilharco et al., 2023; Zhang et al., 2024; Yoshida et al., 2025). A task vector is defined as the parameter difference between a base pre-trained model $\theta_{\mathrm{base}}$ and a fine-tuned model $\theta_{\mathrm{task}}$. By adding or subtracting this vector within the original weight space (so-called “task arithmetic”), a user can edit or remove the corresponding task-specific behavior without further gradient-based training (Ilharco et al., 2023). Moreover, scaling the task vector grants fine-grained control over the strength of the transferred capability. This approach represents a promising direction, as it directly manipulates parameters while avoiding a costly re-optimization of the entire model.
In addition to these computational benefits, prior work has suggested that separating and analyzing task vectors may enhance interpretability (Cerrato et al., 2025). By isolating the weight updates associated with particular subgroups (e.g., racial or gender demographics), one can potentially trace how the model adapts to each subgroup. This feature is appealing for investigating biases arising from unequal representation in training data, as it highlights which groups require larger shifts in weight space. Nevertheless, open questions persist regarding how well this model editing using task-vector preserves or exacerbates fairness. For instance, improving performance for one demographic might degrade outcomes for another, and it is not yet clear how to balance trade-offs with established fairness metrics such as Demographic Parity (DPD) or Equalized Odds (EOD).
To address this gap, we systematically examine how task-vector editing compares to both traditional full-parameter fine-tuning (FFT) and LoRA, and we further explore whether injecting task vectors into an FFT model offers additional control over fairness. Our experiments focus on hate-speech detection on Llama-7B (Touvron et al., 2023) , measured by subgroup-specific accuracy and widely used fairness metrics. Our contributions and findings are summarized as follows:
  • A thorough comparison of four algorithms (FFT, LoRA, model editing using task-vector, and a hybrid approach injecting task vectors into FFT) in terms of their effects on fairness metrics and overall performance (Figure  2)].
  • An analysis showing that task vectors can substantially improve fairness while preserving accuracy, provided that their scalar coefficients are appropriately tuned (Figure 3).
  • Evidence that merging task vectors for underrepresented subgroups with existing models can adjust fairness outcomes without incurring a significant accuracy drop (Figures 4(a), 4(b) and  5(a)).
Through this analysis, we illustrate how task vectors can reduce risks from a fairness perspective while taking advantage of their flexibility and interpretability as a model editing approach. These findings provide a foundation for extending task-vector-based methods to promote fair and responsible operation of large language models.
(a) Gender-based demographic subgroups
(b) Race-based demographic subgroups
Figure 2: LoRA and FFT vs. Task addition with the optimal coefficient for the training accuracy ($\lambda=0.8$ for gender setting and $\lambda=0.5$ for race setting) on group-wise accuracy, demographic parity difference (DPD, lower is fairer), and equalized odds difference (EOD, lower is fairer). Error bars denote the standard error across three seeds. Columns: group-wise accuracy, DPD, EOD. No consistent pattern emerges that task addition necessarily degrades subgroup fairness relative to LoRA or FFT subgroups show improvements or comparable results under task addition, while others show small declines.
2 Preliminaries
In this section, we first provide an overview of the fundamental concept of task vectors and the procedure known as task arithmetic, which applies these vectors to edit model behavior. We then introduce methods for merging multiple task vectors into a single model.
Task arithmetic.
A task vector is defined as the difference in model parameters between a fine-tuned model on a given task and the original base model. Formally, if $\theta_{\text{base}}$ are the pre-trained weights and $\theta_{\text{task}}$ are the weights after fine-tuning on a task, then the task vector is: $\Delta\theta=\theta_{\text{task}}-\theta_{\text{base}}$ (Ilharco et al., 2023).
This vector represents a direction in weight space such that moving the base model’s weights by $\Delta\theta$ steers the model to perform well on that task. In other words, adding $\Delta\theta$ to $\theta_{\text{base}}$ yields a model with improved performance on the target task, without any additional training. Once computed, task vectors can be manipulated through simple arithmetic operations to edit model behavior directly in weight space (Ilharco et al., 2023; Ortiz-Jimenez et al., 2024). Key operations include:
Addition: Given two task vectors $\Delta\theta_{A}$ and $\Delta\theta_{B}$ (for tasks A and B), their sum can be applied to the base model ($\theta_{\text{base}}+\Delta\theta_{A}+\Delta\theta_{B}$) to produce a model that exhibits improved performance on both tasks A and B (Ilharco et al., 2023). This task addition effectively combines knowledge from multiple tasks into one model.
Negation: Using the negative of a task vector, $-\Delta\theta$, one can subtract a task’s influence. For example, applying $\theta_{\text{base}}-\Delta\theta_{A}$ (or equivalently $\theta_{\text{base}}+(-\Delta\theta_{A})$) yields a model with reduced performance on task A—effectively unlearning or forgetting it—while preserving other behaviors (Ilharco et al., 2023). This is useful for removing undesirable skills or biases.
Scalar scaling: Multiplying a task vector by a scalar $\lambda$ adjusts the strength of the edit. For example, using $\theta_{\text{base}}+\lambda\Delta\theta_{A}$ allows partial ($10<\lambda<$ 1) or amplified ($\lambda>1$) application of a task’s effect. This scaling provides fine-grained control over how strongly the task knowledge is injected into the model.
Merging task vectors.
Since task vectors reside in a common weight space, they can be merged by simple addition with tunable scaling. Formally, given a base model $\theta_{0}$ and task vectors $\Delta\theta_{i}$, one can construct a merged model as:
$\theta_{\text{merged}}=\theta_{0}+\sum_{i}\lambda_{i}\,\Delta\theta_{i}\,,$ (1)
where each coefficient $\lambda_{i}$ controls the influence of task $i$. Varying $\lambda_{i}$ thus directly modulates how strongly the $i$-th task’s knowledge is injected, allowing fine-grained blending of capabilities. Indeed, adding multiple task vectors with $\lambda_{i}=1$ endows a model with all those capabilities simultaneously (Ilharco et al., 2023). Optimizing the $\lambda_{i}$ values (i.e., learning an anisotropic scaling for each vector) further improves the composition by balancing contributions and reducing interference between tasks (Zhang et al., 2024).
3 Related Work
Task arithmetic: efficiency and interpretability.
Task vectors offer a computationally efficient framework for editing and analyzing model behavior. Once a task vector is computed—namely, the weight difference between a base model and its fine-tuned variant (Ilharco et al., 2023; Zhang et al., 2024; Yoshida et al., 2025)—no additional training data or retraining is required to transfer or remove task-specific capabilities. By treating each fine-tuning update as a direction in weight space, practitioners can combine or negate these updates through simple addition or subtraction (Ilharco et al., 2023). This modularity not only reduces computational overhead but also enhances interpretability by isolating the contribution of each task.
Beyond modularity, task arithmetic can reveal valuable information about how and where a model adapts to new tasks. Li et al. (2024) show a near-linear relationship between data size and the norm of a task vector, suggesting that over-represented tasks can dominate weight space shifts in multi-task settings. In addition, the orientation of task vectors can indicate synergies or conflicts among tasks (Li et al., 2025), and decomposing these vectors by layer can pinpoint which parts of the model are most affected (Zhang et al., 2024; Gargiulo et al., 2025). Hence, task vectors offer a promising lens for diagnosing training dynamics and identifying potential biases.
Fairness metrics for LLMs.
Fairness in large language models is commonly evaluated using criteria such as Demographic Parity, Equalized Odds, and accuracy parity. Demographic Parity requires similar positive outcome rates across demographic groups, while Equalized Odds demands that true and false positive rates be equivalent. Accuracy parity checks for consistent predictive performance across groups (Fraenkel, 2020; Kennedy et al., 2020a; Pitoura, 2019; Quan et al., 2023). These metrics are broadly used to detect biases and measure whether a model’s behavior disproportionately disadvantages certain populations.
FFT and LoRA under fairness constraints.
Parameter-efficient methods such as LoRA (Hu et al., 2022) address computational bottlenecks by training only a small set of parameters, yet they do not inherently solve fairness issues. In some cases, LoRA yields comparable subgroup performance to full fine-tuning (Ding et al., 2024b), while in others, it fails to mitigate toxic behaviors or biases (Das et al., 2024). The variance in outcomes depends on factors like the rank of the LoRA matrices, the base model’s quality, and the distribution of training data  (Das et al., 2024).
Merging tasks and fairness considerations.
Despite the potential efficiency gains and interpretability offered by task arithmetic, the merging of task vectors for multiple groups can trigger new challenges. For instance, simply summing vectors may lead to “negative transfer,” where updates beneficial to one subgroup degrade performance for another (Ding et al., 2024a; Yu et al., 2020). In highly imbalanced settings, merging models through supervised fine-tuning can also disproportionately favor majority groups while disadvantaging minorities (Cross et al., 2024). Because fairness does not compose additively, interactions among subgroup-specific task vectors can produce unpredictable shifts in metrics like Demographic Parity and Equalized Odds (Gohar et al., 2023).
Consequently, identifying effective ways to adjust task vectors—such as through scalar scaling—remains a key step toward fairness-aware model editing. This work aims to fill that gap by systematically evaluating how these operations influence both fairness and overall model accuracy.
4 Experimental Setup
4.1 Configuration.
Building on the experimental framework established by Ding et al. (2024b), we adopted their evaluation and experimental procedure to assess the fairness implications of LoRA in comparison to FFT. In our work, we extend this analysis by focusing on how task arithmetic compares to both LoRA and FFT in terms of fairness and performance. The detailed experimental setup is provided in Appendix B.
Gender Subgroups Race Subgroups
Men 817 Asian 311
Non-binary 114 Black 1,007
Trans men 178 Latinx 368
Trans unspecified 173 Native American 153
Trans women 148 Middle Eastern 493
Women 2,057 Pacific Islander 138
Other 59 White 580
Other 302
Total 3,546 Total 3,352
Table 1: Data statistics in the gender and race subgroups.
Dataset.
We use a modified version of the Berkeley D-Lab Hate Speech dataset originally introduced by Kennedy et al. (2020a) and adapted by Ding et al. (2024b), the research we are building upon. Our dataset contains a total of 6,898 tweet-sized text snippets annotated for hate speech and categorized by sensitive attributes: Race and Gender, each further divided into fine-grained subgroups (e.g., Women, Non-binary, Men within Gender) as shown in Table 1. We frame hate speech detection as a binary classification task: given a text snippet, the model predicts whether it constitutes hate speech (e.g., hatespeech in the Gender subset may target Non-binary or Trans Women). Each example includes both the hate speech label and one or more protected attribute annotations (e.g., gender = woman, race = Asian). These are used to assess subgroup-level performance and fairness metrics.
This setting supports rigorous fairness analysis due to its rich attribute annotations and real-world relevance. Hate speech detection is a challenging and high-stakes classification problem: it requires models to identify subtle or implicit harm, resolve linguistic ambiguity, and perform robustly across diverse dialects and identity references (Kennedy et al., 2020a). As models increasingly mediate content moderation, ensuring reliable and equitable hate speech detection is essential for safe deployment in real-world systems.
Evaluation metrics.
We evaluate each method on both predictive performance and fairness metrics. Our goal is to understand how scaling or merging task vectors affects these measures.
  • Predictive Performance:
    Accuracy:
    Standard metric for classification tasks, measuring the percentage of correct predictions.
  • Fairness Metrics: We also adopt the metrics used for Ding et al. (2024b) to quantify disparate performance across protected subgroups. These metrics are widely used for fairness research in ML:
    Demographic Parity Difference (DPD):
    Measures the disparity in the model’s positive prediction rates across sensitive attribute groups. A smaller DPD indicates that the model assigns positive outcomes at similar rates across these groups, reflecting a more uniform treatment irrespective of group membership (Agarwal et al., 2018, 2019).
    Equalized Odds Difference (EOD)
    : Measures the disparity in the model’s true and false positive rates across sensitive attribute groups. A smaller EOD indicates that the model’s overall error rate is more balanced across groups (Das et al., 2024).
4.2 Protocol.
We evaluate our methods using a main base model: LLaMA2-7B [^1] LLaMA 2 is licensed under the LLAMA 2 Community License, Copyright (c) Meta Platforms, Inc. All Rights Reserved. See: https://ai.meta.com/llama/license. Our fairness evaluations focus on two sensitive attributes: gender and race, using subgroup-wise metrics mentioned earlier –accuracy, DPD, and EOD.
For FFT, the pretrained model was fine-tuned on the combined training data from all subgroups of the target attribute (gender or race). Evaluation was then performed on the test data from each corresponding subgroup, enabling fine-grained assessment of both performance and fairness.
For LoRA, we followed the same training and evaluation procedure as FFT. The rank of LoRA’s adaptation modules was set to 8, following Ding et al. (2024b).
For task arithmetic, we applied a compositional fine-tuning approach. The training data was partitioned by subgroup (gender or race), and FFT was applied separately to each subgroup’s data to produce fine-tuned models $\theta_{i}$. From these, we computed task vectors $\Delta\theta_{i}$ relative to the base model. These vectors were then merged using the approach described in Eq. (1), with a single, uniform scaling coefficient $\lambda$ applied to all vectors. $\lambda$ served as the sole hyperparameter in the merging process and was tuned on the training data. The evaluation metrics were computed in the same manner as for FFT and LoRA.
Task vector coefficient adjustment.
Building on the task vector merging framework introduced in Eq. (1), we further explore the impact of the scaling coefficient $\lambda$ on fairness outcomes. Specifically, we vary the uniform task vector coefficient $\lambda$ across a broad range (from 0.0 to 1.0 with 0.1 intervals) and evaluate how this adjustment influences subgroup-level fairness metrics, including accuracy, DPD, and EOD.
Impact of worst-performing subgroup task vectors on fairness and performance.
To investigate whether incorporating task vectors from underperforming subgroups can improve fairness without sacrificing overall performance, we first identified the lowest-performing subgroups within each attribute based on the average of DPD and EOD under the FFT setting. We excluded the "others" group from this analysis as it does not reflect the characteristics of any specific subgroup. This selection was informed by both our experimental results and those reported in Ding et al. (2024b), which showed consistent patterns. For gender, the worst-performing subgroups were men and women; for race, they were Asian and Native American. We constructed a new model variant by injecting a worst-performing subgroup task vector worst-performing subgroup task vector into the base fine-tuned model:
$\theta_{\text{new}}=\theta_{\text{SFT}}+\lambda(\theta_{\text{worst-performingsubgroup}}-\theta_{0})$
where $\lambda$ controls the strength of the task vector injection. We varied $\lambda$ from from 0.0 to 1.0 at 0.2 intervals to analyze the effect of this targeted addition on subgroup fairness metrics and overall accuracy.
5 Results
Figure 3: Varying the task arithmetic coefficient $\lambda$ and comparing against FFT (purple dashed) and LoRA (orange dashed) for macro-averaged accuracy (left), demographic parity difference (DPD, center), and equalized odds difference (EOD, right). Evaluated on the gender subset of the data. Higher accuracy is better, while lower DPD and EOD indicate improved fairness. As $\lambda$ changes, task arithmetic maintains competitive accuracy and can reduce fairness gaps relative to the baselines.
Overview.
In Figure 2(a), we compare the performance of FFT, LoRA, and task addition across gender subgroups; Figure 2(b) presents results for race subgroups. For task addition, we selected $\lambda=0.8$ for gender, $\lambda=0.5$ for race, as it achieved the highest average training accuracy across three random seeds within the tested range $\lambda\in[0.0,1.0]$. These visualizations provide a direct comparison of subgroup-wise model behavior.
From the subgroup-level bar plots in Figure 2, we observe that accuracy remains consistently high and comparable across all three adaptation methods, regardless of subgroup.
However, fairness metrics (DPD and EOD) show notable variation across methods. The impact of task addition on fairness is not consistent. Compared to FFT, task addition improved fairness in 5 out of 7 gender subgroups and in 3 out of 8 race subgroups. No single method yielded the best fairness performance across all demographic groups.
Our experimental results closely align with the findings reported in Ding et al. (2024b), particularly regarding the performance of FFT and LoRA on macro-averaged accuracy, DPD, and EOD. This consistency across demographic categories and with prior literature reinforces the robustness of our observations and supports the reliability of our evaluation framework.
5.1 Controlling accuracy and fairness metrics through lambda.
Figure 3 illustrates the overall performance of FFT, LoRA, and task arithmetic as the scaling coefficients for task addition vary from 0.0 to 1.0. We observe how varying the task-arithmetic coefficient $\lambda$ impacts macro-averaged accuracy (left), demographic parity difference (DPD, center), and equalized odds difference (EOD, right) on a gender subset of the data. As $\lambda$ increases from 0.0 to 0.2, we observe a peak in accuracy, but this configuration yields higher DPD and EOD, indicating reduced fairness. Beyond $\lambda=0.3$, accuracy remains competitive compared to FFT and LoRA, while both DPD and EOD progressively decline, suggesting that fairness improves without severely compromising performance. Notably, these task addition curves stay consistently lower than FFT and LoRA in terms of DPD and EOD at higher $\lambda$ values. Overall, this ablation could indicate that tuning $\lambda$ provides a practical mechanism for balancing accuracy and fairness objectives, offering guidelines for practitioners who wish to fine-tune fairness outcomes while maintaining strong predictive performance.
(a) When Men task vector added to the FFT model on the gender subset.
(b) When Women task vector added to the FFT model on the gender subset.
Figure 4: Heatmaps of Accuracy (left), DPD (center), and EOD (right) for gender (top) and race (bottom) subgroups under the baseline FFT model ($\lambda=0.0$) and with increasing $\lambda$ values from 0.2 to 1.0 in 0.2 increments. The task vector for Men was added on the gender subset (top), and the task vector for Women was added on the gender subset (bottom). Darker cells indicate higher values on each metric’s scale; for DPD/EOD, lower values are better.
(a) When Men task vector added to the FFT model on the gender subset.
(b) When Women task vector added to the FFT model on the gender subset.
Figure 5: Impact of injecting both the Men and Women subgroup task vectors into the FFT model on the gender data subset. The plot illustrates how scaling coefficient $\lambda$ reduces DPD and EOD, outperforming the baseline FFT (blue dashed) and LoRA (orange dashed), with negligible impact on macro-averaged accuracy.
5.2 Mixed trends when adding worst-performing subgroup task vectors
To further analyze the effects of subgroup-specific task composition, Figure 4(a)4(b) illustrate heatmaps where the y-axis lists each method or configuration under evaluation: FFT as baseline, followed by task arithmetic with varying scaling coefficients (0.0 to 1.0 with 0.2 intervals). The x-axis represents the subgroups— (e.g., Women, Trans, etc. for Gender). Each cell shows the corresponding performance metric (e.g., macro-averaged accuracy, DPD, or EOD for a given method on a specific subgroup. For these experiments, we added the task vector of the worst-performing subgroups (Women and Men for the gender dataset subset, and Asian, and Native American for the race dataset subset) to the FFT model, as explained earlier.
We generally observe that increasing the scaling coefficient $\lambda$ tends to improve overall accuracy, consistent with the trends observed in Figure 3. However, the impact on fairness metrics (DPD and EOD) is more variable. In the gender-based plots, for example, the Asian subgroup consistently achieves the highest accuracy and lowest DPD/EOD—highlighting a recurring tradeoff where performance gains for one group may exacerbate disparities for others. When the Women task vector is added (Figure 4(b)), accuracy improves for the Trans Women subgroups. However, fairness metrics for subgroups such as Men tend to worsen as the scaling coefficient $\lambda$ increases. In Figure 4(a), injecting the Men task vector improves performance for some subgroups, yet Women consistently show lower accuracy and do not see consistent fairness improvements at higher $\lambda$. Some groups (e.g., Other, Trans Men, Trans Women) begin with relatively poor fairness under FFT and show partial improvements with task vector addition. Still, these improvements are not universal—for example, the Other subgroup often retains high EOD values regardless of $\lambda$. Likewise, Native American accuracy remains mostly unchanged across $\lambda$, while fairness metrics can deteriorate when injecting task vectors for other groups.
Overall, while increasing $\lambda$ can improve both accuracy and fairness for certain subgroups, these effects are not consistent across all configurations.
To visualize these results in more detail, Figure 5(a) shows macro-averaged accuracy, DPD, and EOD for the Men task vector added to the FFT model. The plots illustrate how varying the scaling coefficient $\lambda$ impacts overall performance and fairness, highlighting the effects of subgroup-specific task injection. We can observe in Figure 5(a) that injecting the Men task vector into the FFT model results in a slight accuracy gain and a clear monotonic decrease in both DPD and EOD as $\lambda$ increases—indicating a favorable and consistent improvement in fairness on the gender subset.
However, Figure 5(b) and the additional plots in Figures 11 and 12 in Appendix C.2 show more varied patterns as seen on Figures 4(a) and 4(b). When injecting the Native American task vector (Figure 12), accuracy remains stable while fairness seems to decrease (increased DPD and EOD). Asian (Figure 11) shows the same behavior as injecting the Men task vector (Figure 5(a)), positive increase of fairness metrics as $\lambda$ increases.
Taken together, these results confirm that task vector injection can shift subgroup-wise fairness and performance, but its effects are highly group-specific. While some subgroups (e.g., Men, Asian) exhibit smooth fairness gains, others (e.g., Women) show erratic behavior. This reinforces our earlier conclusion: simple task addition is not a guaranteed to have positive fairness influence, and more targeted strategies are likely needed - yet the scaling coefficient does show some relevant influence in fairness metrics.
6 Conclusion and Limitations
Conclusion.
In this study, we investigated the impact of a task arithmetic approach using task vectors on fairness, in comparison to conventional FFT and LoRA methods. We conducted detailed experiments to assess how the task addition affects prediction accuracy and fairness metrics, including the DPD and EOD across various subgroups. The results indicate that, with appropriate settings of the scalar coefficient $\lambda$, the task arithmetic method can improve DPD and EOD without significantly compromising overall model accuracy. Notably, using low to moderate values of the task vector coefficient effectively reduced prediction bias in minority groups compared to FFT and LoRA.
Furthermore, the task arithmetic framework allows for subgroup-specific evaluation and adjustment of model updates, enhancing interpretability—a key advantage of this method in the context of fairness. This interpretability facilitates the mitigation of excessive bias or adverse effects on particular groups, ultimately enabling more balanced model training.
Limitations.
Despite these promising results, several challenges remain. The effectiveness of task arithmetic depends on dataset characteristics and subgroup distributions, necessitating further investigation into its generalizability across different tasks and domains. Moreover, future work should explore algorithms for automatically optimizing the scalar coefficient $\lambda$ and for balancing trade-offs among multiple subgroups.
In summary, our study demonstrates that task arithmetic using task vectors offers a promising approach for controlling model fairness. Further experimental validation, application to diverse tasks, and developing trade-off optimization methods are essential for improving fairness in broader and more realistic deployment scenarios.
Appendix A Fairness metrics
A.1 Demographic Parity Difference (DPD) (Agarwal et al., 2018, 2019)
DPD measures how varied the model’s rate of positive predictions are across attributes. This metric is calculated as follows:
$M_{\text{DPD}}$ $=\Bigl{|}\Pr[f(X)=1\mid A=1]$
$\quad-\Pr[f(X)=1\mid A=0]\Bigr{|}$ (2)
where $A$ is the sensitive attributes, $f(X)$ is the prediction from the models, and $X$ is the feature vector. The larger the DPD, the greater the difference in prediction outcomes across attributes, indicating greater unfairness in the model predictions.
A.2 Equalized Odds Difference (EOD) (Ding et al., 2024b)
EOD is a metric that measures whether the model exhibits similar predictive performance in terms of true and false positives, regardless of the attribute.
$M_{\text{eod}}=\max\left\{M_{\text{TP}},M_{\text{FP}}\right\}$ (3)
Here, letting $Y$ denote the true label, $M_{TP}$ and $M_{FP}$ are defined as follows:
$M_{\text{TP}}$ $=\Bigl{|}\Pr[f(X)=1\mid Y=1,A=1]$
$\quad-\Pr[f(X)=1\mid Y=1,A=0]\Bigr{|},$ (4)
$M_{\text{FP}}$ $=\Bigl{|}\Pr[f(X)=1\mid Y=0,A=1]$
$\quad-\Pr[f(X)=1\mid Y=0,A=0]\Bigr{|}$ (5)
A.3 Accuracy Parity
Accuracy parity refers to the expectation that a classifier achieves comparable accuracy across different sensitive attribute groups. Formally, accuracy parity is satisfied when the probability of correct classification is equal across groups, i.e.,
$\mathbb{E}(Y=\hat{Y}\mid S=0)=\mathbb{E}(Y=\hat{Y}\mid S=1)$ (6)
This notion of fairness ensures that all subgroups receive equally reliable predictions, and is particularly relevant in applications where consistent model performance across demographics is critical. Unlike statistical parity or equal opportunity, accuracy parity focuses on equal overall correctness rather than specific error types or outcome rates (Quan et al., 2023).
We observed high degree of accuracy parity in both gender and race settings, as the accuracy differences between subgroups are negligible, indicating that the model performs consistently across all groups.
Appendix B Experimental details
B.1 Computational Resources and Software Environment
Hardware and Software:
All experiments presented in this study were performed using computational resources equipped with two NVIDIA H100 GPUs. The experiments leveraged a GPU environment consisting of CUDA 12.1.0, cuDNN 9.0.0, and NCCL 2.20.5.
The experiments were conducted using Python 3.9.18, incorporating several essential Python libraries specifically optimized for deep learning tasks. The primary libraries included PyTorch (version 2.6.0), transformers (version 4.49.0), tokenizers (version 0.21.1), DeepSpeed (version 0.16.4), and Accelerate (version 1.5.2).
The training experiments utilized the DeepSpeed framework with the following key configurations: a gradient accumulation step of 4, optimizer offloaded to the CPU, zero redundancy optimizer at stage 2 (ZeRO-2), and mixed precision training employing FP16 and BF16 for enhanced performance and memory efficiency. All experiments were conducted with a total computational cost of approximately 30 GPU-hours.
Protocol:
We fine-tuned models based on the Llama-7B (Touvron et al., 2023) architecture obtained via HuggingFace repositories.
Each model was trained for 4 epochs, employing a cosine learning rate scheduler with a learning rate of $1\times 10^{-5}$, a warm-up ratio of 0.01, and a weight decay of 0.001. Training utilized a per-device batch size of 2, with an effective batch size of 16 achieved through gradient accumulation. Reproducibility was ensured by setting a random seed of 13, 14, 15 across all experiments.
For Low-Rank Adaptation (LoRA) experiments were conducted with a rank (lora_r) of 8, scaling factor (lora_alpha) of 16, and no dropout.
B.2 Dataset
We use the Berkeley D-Lab hatespeech detection dataset (Kennedy et al., 2020b)  [^2] https://huggingface.co/datasets/ucberkeley-dlab/measuring-hate-speech for our experiments.
The dataset is divided into subgroups based on the following attributes: Race or Ethnicity, Religion, National Origin or Citizenship Status, Gender Identity, Sexual Orientation, Age, and Disability Status. In our study, we use some of these subgroups to evaluate fairness.
Following Das et al. (2024), we binarize the hate speech score associated with each review using a threshold of 0.5 to determine whether the review constitutes hate speech. When multiple annotations exist for the same instance, we obtain one human annotation to avoid duplication.
Appendix C Additional Results
Here, we present results focusing on diverse subgroups, which we could not include in the main paper due to space constraints.
C.1 Comparison of FFT, LoRA, and Task Arithmetic
Figure 6: Boxplots of group-wise accuracy, demographic parity difference (DPD), and equalized odds difference (EOD) for —FFT, LoRA, and task addition with coefficient ($\lambda=0.8$) —evaluated on the gender subset of the data. Higher accuracy is desirable, whereas lower DPD and EOD values indicate improved fairness. Boxplots show medians, interquartile ranges, and variability (with standard error across three seeds). While accuracy is similar across methods, Task Addition generally yields lower DPD and EOD medians than FFT and LoRA, suggesting a better balance between performance and fairness, though overlapping distributions imply these differences are not uniformly significant.
Figure 7: Boxplots of group-wise accuracy, demographic parity difference (DPD), and equalized odds difference (EOD) for —FFT, LoRA, and Task Addition with optimal coefficient ($\lambda=0.5$) —evaluated on the race subset of the data. Higher accuracy is desirable, whereas lower DPD and EOD values indicate improved fairness. Boxplots show medians, interquartile ranges, and variability (with standard error across three seeds).
Figure 8 illustrates the overall performance of FFT, LoRA, and task arithmetic as the scaling for task arithmetic vary from 0.0 to 1.0. Trends observed reinforced results on the gender subset on Figure 3. Overall, $\lambda$ provides a practical mechanism for balancing accuracy and fairness objectives, and similarly there is a peak at $\lambda=0.2$ for highest accuracy, and higher DPD and EOD (less fairness).
Figure 8: On a race-focused subset, we vary task arithmetic’s coefficient $\lambda$ and compare it against FFT (purple dashed) and LoRA (orange dashed). The plots show group-wise accuracy (left), demographic parity difference (DPD, center), and equalized odds difference (EOD, right). Higher accuracy is better, while lower DPD and EOD indicate improved fairness. As $\lambda$ changes, task arithmetic remains competitive in accuracy and can reduce fairness gaps relative to the baselines.
C.2 Subgroup-Specific Task Addition to FFT
We include additional heatmaps that visualize subgroup-wise performance across FFT and varying scaling coefficients for the FFT model injected with a worst-performing subgroup. These supplementary plots, which follow the same setup described earlier, are consistent with the trends observed in Figures 4(a)4(b).
In both gender and race subgroup experiments, increasing the scaling coefficient $\lambda$ generally leads to improved macro-averaged accuracy. However, its impact on fairness metrics—DPD and EOD—is less predictable and varies across subgroups. For instance, some subgroups benefit from improved fairness as their corresponding task vectors are added, while others experience increased disparity, even if accuracy remains stable or improves.
This nuanced behavior reflects a broader pattern: gains in performance for certain subgroups can sometimes come at the expense of fairness for others. Injecting task vectors from worst-performing subgroups does not consistently reduce disparities and, in some cases, can amplify them.
Figure 9: The task vector corresponding to Asian was added to the FFT model on the race data subset. Heatmap of Accuracy (left), DPD (center), and EOD (right) under the baseline (FFT) and increasing $\lambda$ values (0.2 to 1.0). Darker cells indicate higher values in each metric’s scale; for DPD/EOD, lower is better.
Figure 10: The task vector corresponding to Native American was added to the FFT model on the race data subset. Heatmap of Accuracy (left), DPD (center), and EOD (right) under the baseline (FFT) and increasing $\lambda$ values (0.2 to 1.0). Darker cells indicate higher values in each metric’s scale; for DPD/EOD, lower is better.
Figures 125(b) present additional results for the Full+Worst configuration, in which task vectors from the worst-performing subgroups (Native American, Asian, Men, and Women) are added to the FFT model. These plots show macro-averaged accuracy, DPD, and EOD as a function of the scaling coefficient $\lambda$.
Across these figures, we observe mixed effects: while accuracy generally remains stable or improves slightly, fairness outcomes vary by subgroup. In Figure 12, DPD and EOD worsen despite minimal accuracy changes. Meanwhile, Figure 5(b) reveals stable performance with minor fairness improvements, though gains are not consistent across metrics. These results further emphasize that task vector injection alone does not ensure universal fairness improvements and often introduces subgroup-specific trade-offs.
Figure 11: Effect of adding the Asian task vector to the FFT model on the race subset. Accuracy keeps competitive with increasing $\lambda$, and both DPD and EOD decrease consistently.
Figure 12: Results of injecting the Native American task vector into the FFT model. Accuracy shows minimal change across $\lambda$, while DPD and EOD increase (worsen fairness).